# Threat Modeling for AI: The Case for an Asset-Centric Approach

### v1

@@ VI Enterprise RAG: A Secure Retrieval Augmented Generation Architecture
VII Applying Asset-Centric Threat Modeling to Enterprise RAG
To demonstrate the application of asset-centric threat modeling to AI systems, we onboard the Enterprise RAG [25] (Section VI) architecture into this framework. Following the asset-centric methodology we begin by identifying the key AI assets within the system and the potential adversarial capabilities over these assets.
VII-A Asset Identification and Capability Mapping
@@
VII-B Vulnerability Analysis and Threat Contextualization
@@
TABLE II: Asset-Centric Security Analysis of Enterprise RAG Architecture
@@
VIII Related Work
@@ VIII-1 Perspective for Analysis
VIII-2 Analysis Across System Boundaries
@@
Defenders can overcome this challenge through a boundary based approach, enabled by asset-centric analysis. This approach does not require defenders to reason about individual vulnerabilities within systems they lack visibility over. That is, defenders can identify, quantify, and justify their security assumptions about consumed assets at system boundaries – at the point where assets transition across contexts. As described in Section V-B, there are many ways for defenders to justify constraints on their security assumptions. They might, for example, leverage provenance information, the producer’s reputation, or properties about the asset itself. Conversely, producers could leverage an asset-centric perspective to communicate the security guarantees they provide – without needing to open their systems. Even AI-enabling products – such as processors, accelerators and GPUs, or cloud infrastructures – could leverage this perspective to describe the security guarantees they provide for AI assets.
VIII-3 Complementary Analysis
@@
IX Ongoing Work and Next Steps
We have identified, and are pursuing, various opportunities to facilitate the adoption and automation of asset-centric analysis for threat modeling AI. The security analysis in Phase 1 can be facilitated by standardizing the types of AI assets, the range of potential capabilities adversaries can obtain over them (Sections V-A and V-B). The threat analysis in Phase 2 can be facilitated by building and augmenting AI attack knowledge bases with asset-centric information. Defenders can reference such knowledge bases instead of reanalyzing attacks (Section V-C). Finally, threat mapping can be automated using these knowledge bases (Section V-D). These opportunities are, in large part, enabled by standardizing the types of AI assets and mapping out the range of possible capabilities adversaries may obtain over each.
@@ IX Ongoing Work and Next Steps
IX-A Standardizing Assets
As a foundation for facilitating, optimizing, and automating asset-centric analysis, we have enumerated the different types of AI assets and have mapped out the extents to which each can be compromised. Just like for non-AI assets, the range of potential capabilities that an adversary can gain over each type of AI asset can be statically mapped out for each asset type. Today, defenders often denote whether a vulnerability would compromise an asset’s Confidentiality, Integrity, or Availability. These ranges are particularly interesting for AI assets because adversaries can gain previously unseen, nuanced, capabilities over them. One such capability, discussed in the example within Section V-C, is an adversary’s ability to add data to an AI’s training or re-training datasets but lack the ability to read or write the rest of the dataset. Standardization also makes threat models easier to write, interpret, and reuse. Especially for those who are product domain or security experts but not AI security experts. It also facilitates reuse of security assumptions and threat analysis.
IX-B Reusing Security Assumptions
Security assumptions about a consumed asset can be reused across different products, because those assumptions are not specific to the defender’s product. Defenders’ security assumptions about that asset are based on their assumptions about the producer’s systems (Sectio V-B).
@@ IX-B Reusing Security Assumptions
IX-C Automating Threat Analysis
@@
Defenders could programmatically compare each entry in this database – rather than manually – against the potential capabilities they identified in $AF$ (Section V-D).
X Conclusion

### v2

@@ VII Enterprise RAG: A Secure Retrieval Augmented Generation Architecture
VIII Applying Asset-Centric Threat Modeling to Enterprise RAG
To demonstrate the application of asset-centric threat modeling to AI systems, we onboard the Enterprise RAG [24] (Section VII) architecture into this framework. Following the asset-centric methodology we begin by identifying the key AI assets within the system and the potential adversarial capabilities over these assets.
VIII-A Asset Identification and Capability Mapping
@@
VIII-B Vulnerability Analysis and Threat Contextualization
@@
TABLE III: Asset-Centric Security Analysis of Enterprise RAG Architecture
@@
IX Agent-Centric Example: Automated Rust Assistant
To further illustrate our asset-centric methodology in a practical context, we consider Automated Rust Assistant, an AI-powered development agent designed to autonomously write, debug, and optimize Rust code. This example demonstrates how our approach can be applied to agentic AI systems that interact with development ecosystems.
IX-A System Overview
Automatic Rust Assistant is an autonomous AI agent that:
  1. 1.
    Receives high-level programming tasks from users
  2. 2.
    Generates Rust code to fulfill these requirements
  3. 3.
    Executes cargo build commands to compile the code
  4. 4.
    Analyzes compiler errors and warnings
  5. 5.
    Debugs issues using cargo-debugger and VSCode integration
  6. 6.
    Optimizes code based on performance metrics
  7. 7.
    Commits changes to version control systems
The agent has significant capabilities beyond typical AI assistants, including:
  • Direct access to file systems for reading and writing Rust project files
  • Ability to execute shell commands (restricted to cargo ecosystem)
  • Integration with the Debug Adapter Protocol (DAP) used by VSCode to programmatically set breakpoints and receive debugging information
  • Access to version control systems to commit optimized code
IX-B Asset Identification
Following our asset-centric methodology, we identify the key assets in Automated Rust Assistant:
IX-B1 Code Generation Model
  • Pre-trained on Rust codebases and documentation
  • Fine-tuned on high-quality, production Rust code
IX-B2 Error Analysis Model
  • Trained to interpret and resolve Cargo build errors
  • Fine-tuned on real-world debugging sessions
IX-B3 Execution Environment
  • Shell access (limited to cargo commands)
  • File system access (restricted to project directories)
  • Debug Adapter Protocol (DAP) client capabilities for interacting with debuggers
IX-B4 Runtime Behavior Dataset
  • Collection of runtime behavior and performance metrics
  • Used to guide optimization decisions
IX-B5 User Project Codebase
  • Source code files the agent can modify
  • Dependencies and build configurations
IX-B6 Cargo Credentials
  • Potential access to private registry credentials
  • Could enable publishing or installing packages
IX-C Vulnerability Analysis
Using our asset-centric approach, we identify several potential vulnerabilities:
TABLE IV: Asset-Centric Vulnerability Analysis for AutoRust: This table maps key assets to potential adversarial capabilities and corresponding security controls.
Asset Adversarial Capability Security Controls
Code Generation Model Manipulate model to produce malicious or vulnerable code through specially crafted prompts Input validation and sanitization Security-focused model fine-tuning Output scanning for known vulnerability patterns
Debug Adapter Protocol Exploit DAP to access memory contents beyond intended scope Restricted DAP client with limited command scope Runtime monitoring for suspicious memory requests Sandboxed debugging environment
Cargo Dependency Resolution Inject malicious packages through dependency confusion Enforce private registry specifications Checksum verification for all dependencies Dependency lockfiles with integrity checks
Shell Command Interface Execute arbitrary system commands through command injection Restricted command whitelist Execution in isolated container Command argument sanitization
File System Access Read/write sensitive files beyond project scope Path traversal prevention Read-only mountpoints where possible File access auditing
IX-D Threat Modeling Example: Dependency Confusion Attack
Applying our asset-centric methodology, we analyze how a dependency confusion attack could be implemented against AutoRust:
IX-D1 Attack Requirements:
  1. 1.
    User Project Codebase: Write access to Cargo.toml (to modify dependencies)
  2. 2.
    Execution Environment: Ability to execute cargo build (to resolve dependencies)
  3. 3.
    Code Generation Model: Ability to influence code generation to include vulnerable patterns
  4. 4.
    Cargo Credentials: Read access to determine private registry configurations
IX-D2 Attack Implementation:
  1. 1.
    Attacker crafts a prompt that instructs the agent to use an internal dependency
  2. 2.
    Agent updates Cargo.toml with the dependency name but without specifying the private registry
  3. 3.
    When cargo build executes, Cargo first checks the public registry (crates.io)
  4. 4.
    If the attacker has published a malicious package with the same name on the public registry, it gets downloaded instead of the internal package
  5. 5.
    The malicious dependency executes in the context of the project
IX-D3 Asset-Centric Analysis:
  • The vulnerability stems from the agent’s capabilities over multiple assets: write access to project files and ability to execute cargo commands
  • The attack chains these capabilities together to compromise the build process
  • Traditional security would focus on the execution environment, while our approach reveals the full attack path across multiple assets
TABLE V: This table summarizes the Dependency Confusion Attack against AutoRust. Row 3 identifies the asset-centric attack requirements needed to execute this attack.
Dependency Confusion Attack An attack that exploits package resolution mechanisms to inject malicious code into the development environment.
Technique Attacker publishes a malicious package with the same name as an internal dependency to a public registry. When dependency resolution occurs, the malicious package is downloaded instead of the legitimate internal package.
Attack Requirements Adversary must compromise User Project Codebase, Execution Environment, and have knowledge of Cargo Credentials. 1. User Project Codebase: Write access to Cargo.toml to modify dependencies. 2. Execution Environment: Ability to trigger cargo build execution for dependency resolution. 3. Cargo Credentials: Knowledge of internal dependency names without registry specifications.
Impact Malicious code executes within the project’s context with the same privileges as legitimate dependencies, potentially stealing sensitive data or corrupting the development process.
IX-E Mitigations
Based on our methodology, we identify asset-centric mitigations:
IX-E1 Execution Environment Constraints
  • Implement a custom cargo wrapper that always enforces private registry use
  • Add post-execution validation to detect unexpected registry sources
  • Create a secure build sandbox isolated from production environments
IX-E2 Code Generation Model Guardrails
  • Add specific validation for dependency-related operations
  • Train the model to always specify complete registry information
  • Implement pattern recognition for potentially unsafe dependency changes
IX-E3 Integration Boundaries
  • Require human approval for changes to dependency configurations
  • Create checksum verification for critical project files
  • Implement multi-stage verification before executing cargo commands
IX-E4 Debug Adapter Protocol Security
  • Implement a restricted Debug Adapter Protocol client with limited command scope
  • Add runtime monitoring for suspicious memory inspection requests
  • Create validation policies for debugger interactions
  • Develop a sandboxed debugging environment with memory access controls
This example demonstrates how our asset-centric approach provides a comprehensive framework for analyzing complex security challenges in AI agents. By focusing on the assets and the adversarial capabilities over those assets, defenders can identify vulnerabilities that might be missed through traditional security analysis.
Figure 9: Asset-centric view of the AutoRust system architecture showing key components, data flows, and attack vectors. The system consists of an AI Controller that orchestrates multiple subsystems: (1) AI Core Models (including Code Generation and Error Analysis models), (2) Development Environment (with file system access, shell command interface, Cargo build process, and Debug Adapter Protocol integration), and (3) external registries and repositories. Three primary attack vectors are highlighted: prompt injection targeting user input, dependency confusion exploiting registry resolution, and Debug Adapter Protocol exploitation targeting debugging interfaces. Solid arrows represent normal data flows, while dashed red arrows indicate potential attack paths. This architecture illustrates how autonomous AI agents that interact with development tools present unique security challenges that span multiple asset boundaries, requiring comprehensive asset-centric threat modeling.
IX-F Extension: Debug Adapter Protocol Attack
To further illustrate the power of asset-centric analysis, consider a more sophisticated attack that leverages AutoRust’s integration with the Debug Adapter Protocol (DAP) used by debuggers like LLDB and GDB:
IX-F1 Attack Scenario
An attacker with the goal of accessing sensitive information from the debugging session could:
  1. 1.
    Craft a prompt requesting AutoRust to implement a Rust program with legitimate debugging needs
  2. 2.
    Include specific code patterns designed to trigger memory exposure bugs when debugged
  3. 3.
    Manipulate the Debug Adapter Protocol messages exchanged between the DAP client (AutoRust) and the debug adapter
  4. 4.
    Cause the debugger to reveal memory contents beyond the intended scope
  5. 5.
    Extract sensitive data (e.g., encryption keys, authentication tokens) through debug variables
IX-F2 Technical Context
The Debug Adapter Protocol is a JSON-based protocol used by many IDEs (including VSCode) to communicate with language-specific debuggers. AutoRust interacts with this protocol programmatically by:
  • Launching a debug adapter process (e.g., lldb-dap for Rust)
  • Sending JSON-RPC requests to control the debugging session
  • Receiving events and responses containing debugging information
  • Setting breakpoints, examining variables, and evaluating expressions
IX-F3 Asset-Centric Analysis
  • Debug Adapter Protocol: The primary communication channel being manipulated
  • Debug Adapter Process: The target debug engine (LLDB/GDB) processing potentially malicious commands
  • Code Generation Model: Manipulated to create code with specific debugging vulnerabilities
  • Memory Inspector Capabilities: Used to access memory regions outside the intended scope
  • User Project Environment: Contains sensitive data that could be exposed
This more complex attack demonstrates how asset-centric analysis helps defenders reason about cross-boundary attacks that span the AI agent, the debugging protocol, and the underlying memory management system—a vulnerability chain that traditional security analysis might miss, particularly when dealing with AI agents that programmatically control development tools.
X Related Work
@@ X-1 Perspective for Analysis
X-2 Analysis Across System Boundaries
@@
Defenders can overcome this challenge through a boundary based approach, enabled by asset-centric analysis. This approach does not require defenders to reason about individual vulnerabilities within systems they lack visibility over. That is, defenders can identify, quantify, and justify their security assumptions about consumed assets at system boundaries – at the point where assets transition across contexts. As described in Section VI-B, there are many ways for defenders to justify constraints on their security assumptions. They might, for example, leverage provenance information, the producer’s reputation, or properties about the asset itself. Conversely, producers could leverage an asset-centric perspective to communicate the security guarantees they provide – without needing to open their systems. Even AI-enabling products – such as processors, accelerators and GPUs, or cloud infrastructures – could leverage this perspective to describe the security guarantees they provide for AI assets.
X-3 Complementary Analysis
@@
XI Ongoing Work and Next Steps
We have identified, and are pursuing, various opportunities to facilitate the adoption and automation of asset-centric analysis for threat modeling AI. The security analysis in Phase 1 can be facilitated by standardizing the types of AI assets, the range of potential capabilities adversaries can obtain over them (Sections VI-A and VI-B). The threat analysis in Phase 2 can be facilitated by building and augmenting AI attack knowledge bases with asset-centric information. Defenders can reference such knowledge bases instead of reanalyzing attacks (Section VI-C). Finally, threat mapping can be automated using these knowledge bases (Section VI-D). These opportunities are, in large part, enabled by standardizing the types of AI assets and mapping out the range of possible capabilities adversaries may obtain over each.
@@ XI Ongoing Work and Next Steps
XI-A Standardizing Assets
As a foundation for facilitating, optimizing, and automating asset-centric analysis, we have enumerated the different types of AI assets and have mapped out the extents to which each can be compromised. Just like for non-AI assets, the range of potential capabilities that an adversary can gain over each type of AI asset can be statically mapped out for each asset type. Today, defenders often denote whether a vulnerability would compromise an asset’s Confidentiality, Integrity, or Availability. These ranges are particularly interesting for AI assets because adversaries can gain previously unseen, nuanced, capabilities over them. One such capability, discussed in the example within Section VI-C, is an adversary’s ability to add data to an AI’s training or re-training datasets but lack the ability to read or write the rest of the dataset. Standardization also makes threat models easier to write, interpret, and reuse. Especially for those who are product domain or security experts but not AI security experts. It also facilitates reuse of security assumptions and threat analysis.
XI-B Reusing Security Assumptions
Security assumptions about a consumed asset can be reused across different products, because those assumptions are not specific to the defender’s product. Defenders’ security assumptions about that asset are based on their assumptions about the producer’s systems (Sectio VI-B).
@@ XI-B Reusing Security Assumptions
XI-C Automating Threat Analysis
@@
Defenders could programmatically compare each entry in this database – rather than manually – against the potential capabilities they identified in $AF$ (Section VI-D).
XII Conclusion
@@
Appendix A AI Asset Types Cheat-Sheet
Table VI provides an at-a-glance overview of the asset types introduced in Section II. It includes common components, dependencies, and relationships.